Check my work mode : This shows what is correct or incorrect for the work you have completed so far. It does not indicate completion. Return to question 7 Required information [The following information applies to the questions displayed below Duval Co. issues four-year bonds with a $105,000 par value on January 1, 2017, at a price of $100,950. The annual contract 8 points rate is 7%, and interest is paid semiannually on June 30 and December 31. 1. Prepare an amortization table for these bonds. Use the straight-line method of interest amortization. Round your answers to the nearest dollar amount.) Answer is not complete. Semiannual Unamortized Period-End 101/2017 6/30/2017 12/31/2017 6/30/2018 12/31/2018 6/30/2019 12/31/2019 6/30/2020 12/31/2020 Carrying Discount Value 4,050 100,950 506
Answer #1

Semi-annual

Unamortized

Carrying

Period-End

Discount

Value

01/01/2017

$       4,050

$   1,00,950

06/30/2017

$       3,544 *

$   1,01,456

12/31/2017

$       3,038 **

$   1,01,963

06/30/2018

$       2,531

$   1,02,469

12/31/2018

$       2,025

$   1,02,975

06/30/2019

$       1,519

$   1,03,481

12/31/2018

$       1,013

$   1,03,988

06/30/2020

$          506

$   1,04,494

12/31/2020

$              -   

$   1,05,000

*4050-506.25=3544

**3544-506.25=3038 and same is done with others.

Carrying value for every period increase by $506.25

The bond is issued at Discount so the carrying value will increase with every amortization.

Amortization per period= $4050/8= $506.25